OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. 0 = x2 + 5x + 6 0 = x2 + 4x + 4 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 4x + 4 different from y = x2 + 5x + 6?

OpenStudy (anonymous):

does anyone know how to do the questions I posted?

OpenStudy (anonymous):

\[0 = x^2 + 5x + 6 \] factor to write as \[0=(x+3)(x+2)\] set each factor equal to zero to get \[x+3=0\] or \[x+2=0\] solve each get \(x=-3\) or \(x=-2\)

OpenStudy (anonymous):

that was silly the second one is a perfect square \[x^2+4x+4=0\] \[(x+2)^2=0\] \[x=-2\] only one solution to this one

OpenStudy (anonymous):

i will leave the complete sentences up to you, because i am functionally illiterate

OpenStudy (anonymous):

thank you soo much :D